OpenStudy (anonymous):

according to ohms law, what happens to the current as the resistance increases?

OpenStudy (anonymous):

decreases

OpenStudy (anonymous):

\[R=\frac{ E }{C }\]

OpenStudy (amoodarya):

v=R I if v=constant when r increase so I muse decrease like this 10= 5*2 suppose r=5 ,i =2 if r=8 so i=10/8 =5/4=1.25 decreases